# Jet rates in Higgs boson decay at third order in QCD

### v1

@@ Abstract
While the inclusive branching fractions of Higgs bosons to bottom quarks and to gluons are known to fourth order in perturbative QCD [10, 11, 12, 13], predictions for more differential observables such as jet rates or event-shape distributions are restricted to lower orders. Previous precision calculations focused on the decay to bottom quarks, where three-jet production [14] is known to next-to-next-to-leading order (NNLO) in QCD and two-jet production [15] to third order (N3LO). Comprehensive studies of hadronic event shape distributions in both types of decay modes have been performed at next-to-leading order (NLO) in QCD for shape variables related to three-jet [16, 17] and four-jet [18] final states, indicating their potential for discriminating different Higgs boson decay modes [19, 16, 20] in combination with final-state flavour identification [21].
@@
In the first case, the Higgs boson couples to a quark-antiquark pair via Yukawa interaction. Although we consider a non-vanishing bottom-quark Yukawa coupling, light quarks, including the $b$ quark, are treated as massless in the whole calculation. In the second case, the Higgs boson decays into a pair of gluons via a heavy quark loop, which produces an effective vertex in the limit of an infinite heavy-quark mass [22, 23, 24]. The Born-level two-parton decay process corresponds to $\alpha_{\mathrm{s}}^{0}$ for the Yukawa interaction mode and to $\alpha_{\mathrm{s}}^{2}$ for the effective Higgs-to-gluons interaction mode. Perturbative higher-order corrections are defined relative to the coupling order of the Born process throughout. Due to their different chirality structure, the two decay modes do not interfere at any order in perturbation theory.
@@
where $\mu_{R}$ is the renormalization scale,
@@
The coefficients $C^{(n)}_{b\bar{b}}$ and $C^{(n)}_{gg}$ up to N4LO are given for example in [13]. We define the total decay rate to hadrons at NkLO as:
$\Gamma^{(k)}_{H\to\text{had}}=\Gamma^{(k)}_{H\to b\bar{b}}+\Gamma^{(k)}_{H\to gg}.$ (6)
Particles are clustered into jets using a jet algorithm with some pre-defined resolution parameter. We employ the Durham jet algorithm [25, 26, 27, 28, 29], which has been used most widely in jet studies in $e^{+}e^{-}$ annihilation. The resulting Higgs decay rates for different multiplicity then depend on the value of the resolution parameter $y_{\mathrm{cut}}$, and jet multiplicities can be defined by normalizing the Higgs decay rates to the total hadronic Higgs boson decay rate. For both decay categories, we consider a Born-level configuration where the Higgs boson decays to three partons and we compute the relative NNLO correction.
@@
$R^{(k)}_{X}(n,y_{\mathrm{cut}})=\frac{\Gamma_{H\to X}^{(k)}(n,y_{\mathrm{cut}})}{\Gamma_{H\to X}^{(k)}},\,\,\text{with}\,X=gg,\,b\bar{b},\,\text{had},$ (7)
such that the sum of the rates for all possible jet multiplicities at that order sum to unity for any value of $y_{\mathrm{cut}}$,
@@
Since the Higgs boson total decay rate to hadrons is known up to N4LO [10, 11, 12, 13], we can infer the two-jet fraction at N3LO as a function of $y_{\mathrm{cut}}$, $R^{(3)}_{X}(2,y_{\mathrm{cut}})$, from the knowledge of the three-, four- and five-jet fractions, as done in [30] in the context of electron-positron annihilation and associated to $Z$ decays.
@@
The computation of higher-order QCD corrections demands suitable techniques for the removal of infrared (IR) divergences. After renormalization, loop amplitudes contain explicit IR singularities, while the emission of on-shell massless particles leads to the divergent behaviour of real-emission matrix elements in soft and collinear limits. These divergences only cancel when virtual and real corrections are combined in physical predictions.
Figure 2: Normalised three-jet decay rate at LO (green), NLO (blue) and NNLO (red) for Higgs decay to bottom quarks (left), Higgs to gluons (centre), and the weighted sum of the two decay modes (right). The corresponding inclusive decay rate is used as normalisation (see text).
@@
In [43], the detailed validation of all the structures needed for the computation of the NNLO correction to three-jet production in electron-positron annihilation is discussed, and an identical setup is employed here for the Yukawa-mediated mode. The NNLO calculation in the gluonic mode presented in this Letter completes this implementation, now employing the full set of final-state generalized antenna functions [43]. The IR-finite remainders are numerically integrated with the Monte Carlo event generator NNLOjet, suitably adapted for the implementation of generalized antenna functions.
@@
yielding a Higgs vacuum expectation value of $v=\text{246.22}~{}\text{GeV}$. The renormalisation scale is chosen to be $\mu_{\mathrm{R}}=m_{H}$, and we assess theory uncertainties by varying $\mu_{R}$ in the range $[m_{H}/2,2m_{H}]$.
The QCD coupling is set to $\alpha_{\mathrm{s}}(m_{Z})=0.11800$ and its scale evolution is performed with LHAPDF6 [47], yielding $\alpha_{\mathrm{s}}(m_{H})=0.11263$. We take a running $\overline{\mathrm{MS}}$ $b$-quark Yukawa coupling $y_{b}(m_{H})=m_{b}(m_{H})/v=0.011309$. The top-quark mass appears in the two-loop matching coefficient for the infinite top-mass effective field theory [48], and we use the $\overline{\mathrm{MS}}$ top mass $m_{t}(m_{H})=166.48\,\text{GeV}$. Both the $b$-quark Yukawa coupling and the top-quark mass are evolved to different scales according to [49].
@@
In the first two frames of Figure 2, we present the NNLO correction to the normalised three-jet rate for the two different decay modes. To estimate theory uncertainties, the renormalization scale is varied in a correlated manner between numerator and denominator.
@@
We observe that in both channels the NNLO correction is small for large values of $y_{\mathrm{cut}}$, but is significant and negative in regions of lower $y_{\mathrm{cut}}$. The three-jet fraction attains a maximum at a certain value of $y_{\mathrm{cut}}$. Upon inclusion of the NNLO corrections, the height of this three-jet peak decreases and its location is shifted to the right. For the NNLO predictions, the three-jet peak lies at $y_{\mathrm{cut}}\approx 0.002$ and at $y_{\mathrm{cut}}\approx 0.007$ for the Yukawa-induced and the gluonic decay channel respectively, showing a first significant difference between the two modes.
@@
For values of $y_{\mathrm{cut}}$ to the right of the peak, the NNLO result lies within the NLO band, indicating good perturbative convergence, but only a mild reduction of the relative size of scale-variation bands is observed. On the other hand, to the left of the peak, the NLO and NNLO uncertainty bands do not overlap and the two curves exhibit sizeable differences. Indeed, for low $y_{\mathrm{cut}}$ values one expects the emergence of large logarithmic contributions at each order in $\alpha_{\mathrm{s}}$, which spoil the convergence of the perturbative series. In this region, all-order logarithmic resummation is required to obtain sensible predictions for the jet rates. One can consider the location of the three-jet peak as a reasonable estimate for the onset of resummation effects in the respective channels. We observe that for the gluonic decay mode, this happens at values of $y_{\mathrm{cut}}$ about three times as large as the ones for the Yukawa-induced channel.
@@
Figure 3: Fractional jet rates for the production of two (red), three (blue), four (green) and five (yellow) jets at third order in perturbative QCD for Higgs decay to bottom quarks (left) and gluons (centre). The right frame shows the contribution of the two decay modes to the total hadronic two-jet fraction.
@@
In the last frame in Figure 2 we compare the impact of both decay modes on the total three-jet decay rate at NNLO, obtained by the weighted sum of the two contributions, Eq. (6). We observe that the contribution of the gluonic decay mode vanishes for $y_{\mathrm{cut}}\approx 0.001$ and grows to up to $25$% at $y_{\mathrm{cut}}=0.1$. Hard three-jet final states thus offer the highest sensitivity to the gluonic Higgs boson decay mode.
@@
In the right frame of Figure 3, we study the impact of the two decay modes on the total two-jet fraction at N3LO. The contribution of the gluonic decay mode monotonically decreases when smaller values of $y_{\mathrm{cut}}$ are considered. For $y_{\mathrm{cut}}>1/3$ each event is classified as a two-jet event and the Yukawa-induced and the gluonic mode respectively stabilise at $87.4$% and $12.6$% of the total hadronic two-jet rate, in agreement with the contribution of the two modes to the inclusive hadronic branching ratio [50]. By comparing to the right frame of Figure 2, we observe that by selecting hard three-jet final states, the contribution of the gluonic decay mode is doubled compared to a fully exclusive measurement.
In Figure 4 we present the two-jet fractions in hadronic decays of the Higgs at first, second and third order in perturbative QCD.
@@
In this Letter, we performed a fully exclusive calculation of the hadronic decay of a Higgs boson to third perturbative order in the strong coupling. Our calculation accounts for the two dominant decay channels: Yukawa-induced to bottom quarks and heavy-quark-loop-induced to gluons, with the latter not being considered previously at this order. Our results allow us to determine Higgs-boson decay rates into three-jet final states to NNLO and into two-jet final states to N3LO. Our implementation uses the recently proposed generalized antenna functions [43], constructed by means of the designer antenna algorithm [44, 45], for final-state radiation, representing its first application to a new process.
We observe substantial differences between the two channels: for fixed jet resolution $y_{\mathrm{cut}}$ the gluonic mode yields more frequently final states with higher jet multiplicity than the Yukawa-induced process. Consequently, three-jet fraction peaks at larger values of $y_{\mathrm{cut}}$ in the gluonic mode, as expected from the enhancement of radiation patterns off gluons compared to quarks. Our results indicate substantial differences in the phenomenology of hadronic final states between different Higgs boson decay modes, highlighting in particular the enhanced sensitivity on the gluonic decay mode that can be obtained by selecting hard three-jet final states, and provide accurate QCD predictions to exploit these in future precision Higgs studies.

### v2

@@ Abstract
While the inclusive branching fractions of Higgs bosons to bottom quarks and to gluons are known to fourth order in perturbative QCD [10, 11, 12, 13], predictions for more differential observables such as jet rates or event-shape distributions are restricted to lower orders. Previous precision calculations focused on the decay to bottom quarks, where three-jet production [14] is known to next-to-next-to-leading order (NNLO) in QCD and two-jet production [15] to third order (N3LO), with quark mass corrections known up to NNLO [16]. Comprehensive studies of hadronic event-shape distributions in both types of decay modes have been performed at next-to-leading order (NLO) in QCD for shape variables related to three-jet [17, 18] and four-jet [19] final states, indicating their potential for discriminating different Higgs boson decay modes [20, 17, 21] in combination with final-state flavour identification [22].
@@
In the first case, the Higgs boson couples to a quark-antiquark pair via Yukawa interaction. Although we consider non-vanishing bottom-quark Yukawa couplings, light quarks, including the $b$ quark, are treated as massless in the whole calculation. In the second case, the Higgs boson decays into a pair of gluons via a heavy quark loop, which produces an effective vertex in the limit of an infinite heavy-quark mass [23, 24, 25]. The Born-level two-parton decay process corresponds to $\alpha_{\mathrm{s}}^{0}$ for the Yukawa interaction mode and to $\alpha_{\mathrm{s}}^{2}$ for the effective Higgs-to-gluons interaction mode. Perturbative higher-order corrections are defined relative to the coupling order of the Born process throughout. Due to their different chirality structure, the two decay modes do not interfere at any order in perturbation theory in the limit of vanishing bottom quark mass. We explicitly verified that the Born-level interference contributions for massive bottom quarks remain well below one percent for all observables considered here.
@@
where $\mu_{R}$ is the renormalisation scale,
@@
The coefficients $C^{(n)}_{b\bar{b}}$ and $C^{(n)}_{gg}$ up to N4LO are given for example in [13].
Particles are clustered into jets using a jet algorithm with some pre-defined resolution parameter. We employ the Durham jet algorithm [26, 27, 28, 29, 30], which has been used most widely in jet studies in $e^{+}e^{-}$ annihilation. The resulting Higgs decay rates for different multiplicity then depend on the value of the resolution parameter $y_{\mathrm{cut}}$, and jet multiplicities can be defined by normalising the Higgs decay rates to the total hadronic Higgs-boson decay rate. For both decay categories, we consider a Born-level configuration where the Higgs-boson decays to three partons and we compute the relative NNLO correction.
@@
$R^{(k)}_{X}(n,y_{\mathrm{cut}})=\frac{\Gamma_{H\to X}^{(k)}(n,y_{\mathrm{cut}})}{\Gamma_{H\to X}^{(k)}},\,\,\text{with}\,X=gg,\,b\bar{b},\,$ (6)
such that the rates for all possible jet multiplicities at that order sum to unity for any value of $y_{\mathrm{cut}}$,
@@
Since the Higgs boson total decay rate to hadrons is known up to N4LO [10, 11, 12, 13] in either mode, we can infer the two-jet fraction at N3LO as a function of $y_{\mathrm{cut}}$, $R^{(3)}_{X}(2,y_{\mathrm{cut}})$, from the knowledge of the three-, four- and five-jet fractions, as done in [31] in the context of electron-positron annihilation and associated to $Z$ decays.
@@
The computation of higher-order QCD corrections demands suitable techniques for the removal of infrared (IR) divergences. After renormalisation, loop amplitudes contain explicit IR singularities, while the emission of on-shell massless particles leads to the divergent behaviour of real-emission matrix elements in soft and collinear limits. These divergences only cancel when virtual and real corrections are combined in physical predictions.
Figure 2: Normalised three-jet decay rate at LO (green), NLO (blue) and NNLO (red) for Higgs decay to bottom quarks (left), Higgs to gluons (centre), and the weighted sum of the different decay modes (right). The corresponding inclusive decay rate is used as normalisation (see text).
@@
In [44], the detailed validation of all the structures needed for the computation of the NNLO correction to three-jet production in electron-positron annihilation is discussed, and an identical setup is employed here for the Yukawa-mediated mode. The NNLO calculation in the gluonic mode presented in this Letter completes this implementation, now employing the full set of final-state generalised antenna functions [44]. The IR-finite remainders are numerically integrated with the Monte Carlo event generator NNLOjet, suitably adapted for the implementation of generalised antenna functions.
@@
yielding a Higgs vacuum expectation value of $v=\text{246.22}~{}\text{GeV}$. The renormalisation scale is chosen to be $\mu_{\mathrm{R}}=m_{H}$, and we assess theory uncertainties by varying $\mu_{R}$ in the range $[m_{H}/2,2m_{H}]$. The QCD coupling is set to $\alpha_{\mathrm{s}}(m_{Z})=0.11800$ and its scale evolution is performed with LHAPDF6 [48], yielding $\alpha_{\mathrm{s}}(m_{H})=0.11263$.
We specifically focus on the decay to bottom quarks and gluons to expose their differences, but also consider the combination of the two modes. To obtain phenomenologically viable results for the combination of the different decay modes, we also include the Yukawa decay to charm quarks, which is identical to the bottom Yukawa decay up to its overall normalisation. Moreover, we rescale $\lambda_{0}^{2}(\mu_{R})$ in (5) to account for the exact top, bottom and charm mass dependence of the one-loop Higgs-gluon-gluon vertex [49], including interference contributions. The combined result is normalized with respect to the total decay rate to hadrons at NkLO:
$\Gamma^{(k)}_{H\to\text{had}}=\Gamma^{(k)}_{H\to b\bar{b}}+\Gamma^{(k)}_{H\to c\bar{c}}+\Gamma^{(k)}_{H\to gg}.$ (8)
We take running $\overline{\mathrm{MS}}$ bottom and charm quark masses and Yukawa couplings $y_{b}(m_{H})=m_{b}(m_{H})/v=0.011309$, $y_{c}(m_{H})=m_{c}(m_{H})/v=0.0024629$ and use the $\overline{\mathrm{MS}}$ top mass $m_{t}(m_{H})=166.48\,\text{GeV}$. The Yukawa couplings and the top-quark mass are evolved to different scales according to [50].
@@
In the first two frames of Figure 2, we present the NNLO correction to the normalised three-jet rate for the two different decay modes. To estimate theory uncertainties, the renormalisation scale is varied in a correlated manner between numerator and denominator.
@@
We observe that in both channels the NNLO correction is small for large values of $y_{\mathrm{cut}}$, but is significant and negative in regions of lower $y_{\mathrm{cut}}$. The three-jet fraction attains a maximum at a certain value of $y_{\mathrm{cut}}$. Upon inclusion of the NNLO corrections, the height of this three-jet peak decreases and its location is shifted to the right. For the NNLO predictions, the three-jet peak lies at $y_{\mathrm{cut}}\approx 0.002$ and at $y_{\mathrm{cut}}\approx 0.007$ for the $b$-quark and the gluonic decay channel respectively, showing a first significant difference between the two modes.
@@
For values of $y_{\mathrm{cut}}$ to the right of the peak, the NNLO result lies within the NLO band, indicating good perturbative convergence, and the relative size of the theory uncertainties is halved. On the other hand, to the left of the peak, the NLO and NNLO uncertainty bands do not overlap and the two curves exhibit sizeable differences. Indeed, for low $y_{\mathrm{cut}}$ values one expects the emergence of large logarithmic contributions at each order in $\alpha_{\mathrm{s}}$, which spoil the convergence of the perturbative series. In this region, all-order logarithmic resummation is required to obtain sensible predictions for the jet rates. One can consider the location of the three-jet peak as a reasonable estimate for the onset of resummation effects in the respective channels. We observe that for the gluonic decay mode, this happens at values of $y_{\mathrm{cut}}$ about three times as large as the ones for the Yukawa-induced channel.
@@
Figure 3: Fractional jet rates for the production of two (red), three (blue), four (green) and five (yellow) jets at third order in perturbative QCD for Higgs decay to bottom quarks (left) and gluons (centre). The right frame shows the contribution of the different decay modes to the total hadronic two-jet fraction.
@@
In the last frame in Figure 2 we compare the impact of the different decay modes on the total three-jet decay rate at NNLO, obtained by the weighted sum of the bottom, charm, and gluonic contributions normalized to (8). We observe that the contribution of the gluonic decay mode vanishes for $y_{\mathrm{cut}}\approx 0.001$ and grows to up to $25$% at $y_{\mathrm{cut}}=0.1$. Hard three-jet final states thus offer the highest sensitivity to the gluonic Higgs boson decay mode.
@@
In the right frame of Figure 3, we study the impact of the different decay modes on the total two-jet fraction at N3LO. The contribution of the gluonic decay mode monotonically decreases when smaller values of $y_{\mathrm{cut}}$ are considered. For $y_{\mathrm{cut}}>1/3$ each event is classified as a two-jet event and the the different decay modes respectively stabilise at $84.7$% (bottom Yukawa), $11.3$% (gluonic) and $4.0$% (charm Yukawa) of the total hadronic two-jet rate, in agreement with their contribution to the inclusive hadronic branching ratio [51]. By comparing to the right frame of Figure 2, we observe that by selecting hard three-jet final states, the contribution of the gluonic decay mode is doubled compared to a fully exclusive measurement.
In Figure 4 we present the two-jet fractions in hadronic decays of the Higgs at first, second, and third order in perturbative QCD.
@@
In this Letter, we performed a fully exclusive calculation of the hadronic decay of a Higgs boson to third perturbative order in the strong coupling. Our calculation accounts for the two dominant types of decay categories: Yukawa-induced to bottom or charm quarks and heavy-quark-loop-induced to gluons, with the latter not being considered previously at this order. Our results allow us to determine Higgs-boson decay rates into three-jet final states to NNLO and into two-jet final states to N3LO. Our implementation uses the recently proposed generalised antenna functions [44], constructed by means of the designer antenna algorithm [45, 46], for final-state radiation, representing its first application to a new process.
We observe substantial differences between the two channels: for fixed jet resolution $y_{\mathrm{cut}}$ the gluonic mode yields more frequently final states with higher jet multiplicity than the Yukawa-induced process. Consequently, three-jet fraction peaks at larger values of $y_{\mathrm{cut}}$ in the gluonic mode, as expected from the enhancement of radiation patterns off gluons compared to quarks. Our results indicate substantial differences in the phenomenology of hadronic final states between different types of decay categories, highlighting in particular the enhanced sensitivity on the gluonic decay mode that can be obtained by selecting hard three-jet final states, and provide accurate QCD predictions to exploit these in future precision Higgs studies.